OpenStudy (anonymous):

Evaluate the following expression using the values given: Find -a^2 - 3b^3 + c^2 + 2b^3 - c^2 if a = 3, b = 2, and c = -3

OpenStudy (anonymous):

ik lol but i got lost after i did that step

OpenStudy (anonymous):

-(3)2 - 3(2)3 + (-3)2 + 2(2)3 - (-3)2 is the equation after you substitute the values for its corresponding letter

OpenStudy (anonymous):

oh wait hang on lol

OpenStudy (anonymous):

sorry look at the q again pls :/

OpenStudy (astrophysics):

\[-(3)^2-3(2)^3+(-3)^2+2(2)^2 = ?\]

OpenStudy (anonymous):

yes

OpenStudy (astrophysics):

Ok what did you get

OpenStudy (anonymous):

i got lost in that step :/

OpenStudy (astrophysics):

What step, all we have to do is plug in the values then evaluate

OpenStudy (anonymous):

the evaluating part

OpenStudy (astrophysics):

\[(3)^2 = (3)(3),~~~~(2)^3 = (2)(2)(2)\]

OpenStudy (astrophysics):

That's what it means

OpenStudy (anonymous):

ik

OpenStudy (astrophysics):

Ok so you can write it out like that, try evaluating it now

OpenStudy (anonymous):

ok